# {textblock}

### v1

@@ Introduction
Table 1 provides information on the recognition accuracy of each model under benign conditions (i.e., with clean data) and attack conditions (i.e., with 5% poisoned data). It can be seen that the accuracy under attack conditions is almost similar for all models in terms of performance, except RNN with Attention, which has a 100% rate. (Koffas et al. 2022) (as the attack fits perfectly into the time-dependency level).
Table 1: Performance comparison of backdoored models
@@
CNN 97.31% 99.75%
VGG16 99.06% 99.55%
CNN-LSTM 96.67% 99.25%
RNN with Attention 96.06% 100%
CNN-RNN 94.63% 99.0%
LSTM 74.12% 99.0%
@@
The proposed data poisoning technique (Figure 2) is based on the creation of a function that, by inserting sounds into clean audio data, creates dynamic triggers. Imperceptibility checks are included in this dynamic audio data poisoning to ensure that the triggers created cannot be distinguished by human listeners. This allows minor modifications to be made without producing audible artifacts. As part of the poisoning process, a backdoor subtly modifies the original audio samples using the dynamic trigger. The specified target label of the misclassification allows for flexibility in choosing the desired result.
Figure 3 shows the results obtained by DynamicTrigger for the benign model (BA) after poisoning the dataset. The poisoned data is shuffled. To construct new data sets, we mix the samples from the clean training set with those from the poisoned training set. To do this, we concatenate parts of the original training and test datasets, resulting in mixed samples for inputs and labels. Finally, the BA model(s) are cloned and generated in such a way that they can be trained on this mixed data (to finally obtain the poisoned ASR model(s) Table 1).
Figure 3: Top plots show three separate clean spectrograms and bottom plots their poisoned (backdoored) counterparts with decisions made by the CNN-LSTM model ( Table1).
As can be seen in (Figure 3), the effectiveness of the proposed method is demonstrated on a set of clean audio samples ("clean audio clip") with a specified target label. The backdoor process successfully introduces imperceptible triggers when training the DNN model(s), resulting in poisoned audio samples ("DynamicTrigger audio clip"). Model predictions on the poisoned samples ("DynamicTrigger audio clip") show the misclassification of the desired target.
Characterizing the effectiveness of DynamicTrigger.
Two methods (Figures 4, 5 and 6 ) are used to assess the risk of DynamicTrigger backdoor attacks: Activation Defense (Nicolae et al. 2018) and a dimensionality reduction technique (T-SNE PCA)(Soremekun, Udeshi, and Chattopadhyay 2023).
@@
Activation Defense without reclassification (see Figure 5), this approach relies on anomaly detection algorithms such as DBSCAN (Alfoudi et al. 2022) and is not always as effective at eliminating the backdoor’s influence. Because it does not contain explicit information on the location of the backdoor in the DNN network.
@@
In Figure 5, Activation Defense detected DynamicTrigger during detection without reclassification and was able to detect backdoored clusters, but the problem with this method is that it does not suppress the attack and incorporates many false positives.
Activation Defense with reclassification (see Figure 6), this approach eliminates the backdoor’s influence on the network more strongly and permanently, as identified neurons or layers are directly eliminated or modified.
@@
In Figure 6, Activation Defense did not detect DynamicTrigger during detection with reclassification and was unable to detect backdoored clusters.
@@
Figures 5 and 6 show that the "Activation Defense" method (for further explanation of this defense method, see the reference article (Nicolae et al. 2018)) on the audio backdoor attack has great difficulty in accurately detecting DynamicTrigger, leading to the conclusion that this dynamic backdoor attack is stealth (for more exhaustive results, see code ***code).
conclusion.
@@
This paper presents an efficient dynamic backdoor attack approach for voice command recognition by injecting triggers into the auditory speech spectrum. The experimental results illustrate that DynamicTrigger can achieve a sufficiently high ASR while remaining sufficiently stealthy. Additionally, DynamicTrigger proves effective in resisting standard defense methods. Existing backdoor attacks mostly concentrate on classification tasks, while triggers designed for voice command recognition also demonstrate efficacy in other such tasks, including speaker recognition. This article aims to draw the attention of financial services to the adoption of such authentication mechanisms and to encourage them to implement, in addition to these authentication techniques, other means of ensuring that their voice recognition authentication systems are preserved and protected against backdoor attacks. Our aim for the future is to expand DynamicTrigger to wider automatic recognition tasks.
@@ Acknowledgments.
Appendix (additional results).
Dataset.
To extend the results of our DynamicTrigger model of dynamic backdoor attack for generalization purposes, we use the TIMIT corpusdocumentation of read speech is intended to provide speech data for acoustic and phonetic studies, as well as for the development and evaluation of automatic speech recognition systems. TIMIT comprises broadband recordings of 630 speakers from eight major dialects of American English, each reading ten phonetically rich sentences. The TIMIT corpus comprises time-aligned orthographic, phonetic, and verbal transcriptions, along with a 16-bit, 16 kHz speech waveform file for each utterance.
Table 2: Performance comparison of backdoored models
Models Benign Accuracy (BA) Attack Success Rate (ASR)
CNN 97.63% 100%
VGG16 99.06% 99.55%
CNN-LSTM 87.81% 100%
RNN with Attention 96.06% 100%
CNN-RNN 87.31% 100%
  • 1
    630 speakers ; DARPA TIMIT Acoustic-phonetic continuous.
Figure 7: Poisoning of TIMIT database data through successful activation of the "backdoor" tag. The top graphs show three distinct clean spectrograms (for each respective speaker with its Unique ID (label)) and the bottom graphs their respective poisoned equivalents.
Figure 8: Backdoor attack (DynamicTrigger) on the TIMIT database (already poisoned at this stage) through successful activation of the "3" label. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned (backdoored) equivalents (by DynamicTrigger), with decisions made by the CNN-LSTM model (Table 2).
Speaker verification
Objective :
speaker recognition, on DNN models trained with the TIMIT database poisoned by the DynamicTrigger attack, we apply the speaker verifier(s), ECAPA-TDNN (Ravanelli et al. 2021)(see Figure 9) of HugginFace HugginFace speaker recognition and Nvidia’s Speakernet (Koluguri, Park, and Ginsburg 2022) (see Figure 10) to detect irregularities in the final audio data (poisoned) obtained by the DNN model(s) trained on this data in comparison with the clean data.
Figure 9: Speaker verification ECAPA-TDNN.
Figure 10: Speaker verification NeMo Nvidia.

### v3

@@ VI-A Backdoor Attack Performance and discussion:
Table II provides information on the recognition accuracy of each model under benign (i.e., with clean data) and attack conditions (i.e., with 0.2% poisoned data). It can be seen that the accuracy under attack conditions is almost similar for all models in terms of performance, except for RNN with Attention, which has a 99% rate. [35] (because the attack fits perfectly into the time-dependency level).
Table II: Performance comparison of backdoored models
@@
CNN 97.31% 100%
VGG16 99.06% 100%
CNN-LSTM 96.67% 100%
RNN with Attention 96.06% 99.0%
CNN-RNN 94.63% 100.0%
LSTM 74.12% 100.0%
ResNet-34 76.12% 100.0%
ResNet-50 78.12% 100.0%
@@
The proposed data poisoning technique (Figure 3) is based on the creation of a function that creates dynamic triggers, by inserting sounds into clean audio data. Imperceptibility temporal-distributed trigger [22] checks were included in this dynamic audio data poisoning to ensure that the triggers created could not be distinguished by human listeners. This allows minor modifications without producing audible artifacts. As part of the poisoning process, the backdoor subtly modifies the original audio samples using a dynamic trigger.
@@
Figure 4 shows the results obtained by DynamicTrigger for the benign model (BA) after poisoning the dataset. The poisoned data were then shuffled. To construct new data sets, we mixed samples from the clean training set with those from the poisoned training set. To do this, we concatenated parts of the original training and test datasets, resulting in mixed samples for inputs and labels. Finally, the BA model(s) are cloned and generated in such a way that they can be trained on the mixed data (to finally obtain the poisoned ASR model(s) Table in II).
Figure 4: Top plots show three separate clean spectrograms and bottom plots their poisoned (backdoored) counterparts with decisions made by the CNN-LSTM model ( TableII).
As shown in (Figure 4), the effectiveness of the proposed method is demonstrated on a set of clean audio samples (”clean audio clip”) with a specified target label. The backdoor process successfully introduceds imperceptible triggers when training the DNN model(s), resulting in poisoned audio samples (”DynamicTrigger audio clip”). Model predictions of the poisoned samples (”DynamicTrigger audio clip”) show misclassification of the desired target.
VI-B Characterizing the effectiveness of DynamicTrigger.
Two methods (Figures 5, 6 and 7) were used to assess the risk of DynamicTrigger backdoor attacks: Activation Defense [29] and the dimensionality reduction technique (T-SNE PCA)[30].
@@
Activation Defense without reclassification (see Figure 6), this approach relies on anomaly detection algorithms such as DBSCAN [60] and is not always as effective at eliminating the influence of the backdoor. This is because it does not contain explicit information regarding the location of the backdoor in the DNN network.
@@
In Figure 6, the Activation Defense detected DynamicTrigger during detection without reclassification and was able to detect backdoored clusters, However the problem with this method is that it does not suppress the attack and incorporates many false positives. Activation Defense with reclassification (see Figure 7), this approach eliminates the backdoor’s influence on the network more strongly and permanently, as identified neurons or layers are directly eliminated or modified.
@@
In Figure 7, the Activation Defense did not detect DynamicTrigger during detection with reclassification and was unable to detect backdoored clusters. Figures 6 and 7 show that the ”Activation Defense” method (for further explanation of this defense method, see the reference article [29]) on the audio backdoor attack has great difficulty in accurately detecting DynamicTrigger, which lead to the conclusion that this dynamic backdoor attack is stealth (for more exhaustive results, see code  [^4] ART.1.18 IBM).
VII Ablation Study.
The Impact of Audio Stylistic Transformations (TranStyBack) on DynamicTrigger Backdoor Attack.
To understand the impact of audio stylistic transformations (TranStyBack) [40] on our DynamicTrigger backdoor attack in the audio domain, we’re developing a new algorithm 2 (for more details, see this link [^5] ART.1.18 IBM) to which we associate DynamciTrigger by incorporating stylistic transformations (TranStyBack, Table III). Backdoor attacks that capitalize on stylistic changes (algorithm 2) usually involve a set of diverse techniques. Their goal is to embed harmful triggers or backdoor functionality that can be reliably controlled by an attacker in parameters, such as the model’s weights or architecture. While conventional backdoor attacks usually prepare a template to be the trigger, stylistic modifications enable something different. For instance, the attacker might create an input that reliably results in a particular kind of dynamic shift in the behavior of the compromised model. One could also view these methods as a kind of ad hoc data poisoning attack, whereby ”trojanized” data is used to train a regular model. We can therefore state the following :
Adversarial risk is defined as:
$R_{\text{adv}}(f,D)=\mathbb{E}_{(x,y)\sim D}\left[\max_{x^{\prime}\in B(x,\epsilon)}\ell(f(x^{\prime}),y)\right],$
where $B(x,\epsilon)$ represents the ball around $x$ with radius $\epsilon$, indicating the space of adversarial.
  1. 1.
    Generalization Bound [61]: A measure of how well a model generalizes [62] to unseen data, often quantified using bounds like Hoeffding’s inequality [63]
    $\Pr\left(\left|\hat{L}_{n}(h)-L_{P}(h)\right|\geq t\right)\leq e^{-\frac{2nt^{2}}{\sigma^{2}}}$
    where $\hat{L}_{n}(h)$ is the empirical error rate, $L_{P}(h)$ is the true error rate, $t$ is the margin of error, and $\sigma^{2}$ is the variance of the loss function. or VC dimension [64]
    $\Pr_{S\sim P^{n}}\left(\exists h\in\mathcal{H}:\forall x\in S,h(x)\neq y\right)\leq\frac{d\cdot|S|}{2^{d}}$
    where $S$ is a set of $n$ points, $P$ is the underlying distribution, and $d$ is the VC dimension of $\mathcal{H}$.
    $\mathcal{R}(f)\leq\mathcal{R}_{\text{emp}}(f)+\sqrt{\frac{\log(n)}{2n}}+\sqrt{\frac{\log(m)}{2m}},$
    where $\mathcal{R}(f)$ is the true error rate of the model, $\mathcal{R}_{\text{emp}}(f)$ is the empirical error rate, $n$ is the number of training samples, and $m$ is the number of classes or labels.
  2. 2.
    Robust Optimization: Minimizing the worst-case loss over a perturbed input space, leading to formulations like:
    $\min_{f}\max_{x^{\prime}\in B(x,\epsilon)}\ell(f(x^{\prime}),y).$
  3. 3.
    Differential Privacy [65]: Ensuring that small changes in the dataset result in minimal changes in the model output, formalized as:
    $\Pr[\mathcal{O}(D^{\prime})-\mathcal{O}(D)>\epsilon]\leq\delta,$
    where $\mathcal{O}$ is an observable, $D^{\prime}$ is a neighboring dataset, and $\epsilon$ and $\delta$ are parameters controlling privacy.
  4. 4.
    Game Theory [66],[67],[68],[69]: Modeling interactions between the model and adversaries as games, leading to strategies like minimax regret:
    $\min_{f}\max_{x^{\prime}\in B(x,\epsilon)}\ell(f(x^{\prime}),y)-\min_{f}\ell(f(x),y).$
  5. 5.
    Rademacher Complexity [70],[71]: For a function class $\mathcal{F}$ and a set of $n$ points $S$, the Rademacher complexity is defined as:
    $\mathcal{R}_{n}(\mathcal{F})=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\sum_{i=1}^{n}\sigma_{i}f(x_{i})\right]$
    where $\sigma$ is a sequence of independent random variables taking values in $\{-1,1\}$, and $f(x_{i})$ is the prediction of $f$ at point $x_{i}$.
VII-A DynamicTrigger Backdoor Attack Using Stylistic Transformation.
VII-A1 Dataset.
To extend the results of our DynamicTrigger model of dynamic backdoor attack for generalization purposes, we use the TIMIT corpus666documentation of read speech is intended to provide speech data for acoustic and phonetic studies, as well as for the development and evaluation of automatic speech recognition systems. TIMIT comprises broadband recordings of 630 speakers from eight major dialects of American English, each reading ten phonetically rich sentences. The TIMIT corpus comprises time-aligned orthographic, phonetic, and verbal transcriptions, along with a 16-bit, 16 kHz speech waveform file for each utterance. On the following transformers [^7] Open ASR Leaderboard (Distil-Whisper, Whisper, MMS, Wav2Vec2, Hubert, Wav2Vec2-BERT). The experimental conditions were the same as those described in the article under V-D.
Table III: Stylistic triggers employed in our experiments.
Style Effect Description
0 PitchShift(S, 40) shifts the pitch of the audio signal by 20 semitones.
1 Distortion(S, 80 dB) adds distortion to the audio signal.
2 Chorus(S, 40 ms, 15) chorus effect by add a delayed version of the signal.
3 Chorus(Distortion(PitchShift(S, 70), 80 dB), 8 ms, 5) creates a chorus effect and adds a delayed and distorted version of the signal.
4 Reverb(Distortion(Chorus(S, 25 ms, 0.25), 90 dB)) creates a reverb effect by simulating the reflections of sound in a room.
5 Phaser(Ladder(Gain(S, 25 dB))) creates a phaser effect and adds a delayed and modulated version of the signal.
Backdoor attack : TranStyBack
Algorithm 2 : Stylistic backdoor attack audio
1:clean audio samples ($A_{c}$), Target label ($T$), Effects ($E$)
2:Poisoned audio samples ($A_{p}$), Poisoned labels ($L_{p}$)
3:procedure StylisticBackdoorAttack($A_{c},T,E$)
4:    $A_{p}\leftarrow\textsc{InsertTrigger}(A_{c})$ Insert trigger audio into clean audio (clapping , samplerate=16khz )
5:    $A_{p}\leftarrow\textsc{ApplyEffects}(A_{p},E)$ Apply stylistic effects (see Table III) to audio
6:    $L_{p}\leftarrow\textsc{AssignLabels}(T)$ Assign poisoned labels
7:    return $A_{p},L_{p}$
8:procedure InsertTrigger($A_{c}$)
9:    Read trigger audio ($T_{\text{audio}}$) from backdoor trigger path Read clapping sound as trigger audio
10:    $A_{p}\leftarrow A_{c}+T_{\text{audio}}$ Concatenate trigger audio with clean audio
11:    return $A_{p}$
12:procedure ApplyEffects($A,E$)
13:    for each effect $e\in E$ do Apply six different audio effects
14:       $A\leftarrow\text{ApplyEffect}(A,e)$
15:    return $A$
16:procedure AssignLabels($T$)
17:    $L_{p}\leftarrow\text{Assign target label(s)}(T)$ Assign target label(s) to poisoned audio
18:    return $L_{p}$
TranStyBack (algorithm 2) focuses on the implementation of an audio backdoor attack using stylistic transformations. Thus, our study explores the possibility of executing such an attack using digital musical effects. For this, we use two frameworks, namely, audiomentations and Pedalboard, to implement six styles combining effects, such as PitchShift, Distortion, Chorus, Reverb, Gain, Ladderfilter, and Phaser. For an overview of the effects considered, see Table III.
Table IV: Performance comparison of backdoored models.
Models Benign Accuracy (BA) Attack Success Rate (ASR)
Whisper 97.63% 100%
MMS 99.06% 100%
Distil-Whisper 87.81% 100%
Wav2Vec2 96.06% 100%
Hubert 87.31% 100%
Wav2Vec2-BERT 74.12% 100%
  • 1
    630 speakers ; DARPA TIMIT Acoustic-phonetic continuous.
  • 2
    These pre-trained models are available on (Open ASR Leaderboard).
Figure 8: Backdoor attack (TranStyBack) on the TIMIT database (already poisoned at this stage) through successful activation of the ”3” label. The top graphs show three distinct clean spectrograms (for each respective speaker with its unique ID (label)), and the bottom graphs show their respective poisoned (backdoored) equivalents (by TranStyBack), with decisions made by the Wav2Vec2-BERT model (Table IV).
VII-A2 Signal Fidelity Analysis : Effect of the Poisoning
Signal Fidelity: To evaluate the fidelity (Figure 9) of the audio stylistic transformations (for three samples), we employed signal-based metrics such as Total Harmonic Distortion (THD) [72],[73],[73],[74], and Signal-to-Noise Ratio (SNR)) [75],[76],[77],[78]. These metrics measure the quality and distortion introduced by stylistic transformations.
Figure 9: Signal Fidelity
SNR $=10\log_{10}\left(\frac{{\|x_{\text{poisoned\_trans}}\|^{2}}}{{\|x_{\text{clean}}-x_{\text{poisoned\_trans}}\|^{2}}}\right)$
THD $=\frac{{\sqrt{{\sum_{n=2}^{N_{\text{freq}}}\|X_{\text{harmonic}}(n)\|^{2}}}}}{{\|X_{\text{fundamental}}\|}}$
where, $x_{\text{poisoned\_trans}}$ represents the transformed poisoned audio signal, $x_{\text{clean}}$ represents the original clean audio signal, $X_{\text{harmonic}}(n)$ represents the $n$-th harmonic component of the audio signal, $X_{\text{fundamental}}$ represents the fundamental component of the audio signal, and $N_{\text{freq}}$ represents the number of frequency components. We utilize the techniques of dimensionality reduction [79], as well as signal fidelity analysis, in an effort to comprehend and detect the attack.
VII-B Dimensionality reduction techniques
Dimensionality reduction techniques such as Locally Linear Embedding (LLE) [80],[81],[82], t-SNE [83],[84],[85],[86] , Spectral Embedding [87, 88] , Isomap [89, 90] and UMAP [91],[92],[93], are used to analyze audio backdoor attacks by transforming high-dimensional audio data into lower-dimensional representations, which provides insight into the underlying structure and relationships within the data, allowing visualization and identification of patterns or anomalies related to the attack (TranStyBack); which can then be displayed to visualize the distribution of output representations (see Figure 10).
Figure 10: Manifold Embedding stylistics backdoor
To demonstrate the effects of audio backdoor attacks that rely on stylistic methods, we use multi-dimensional integration in Figure 10 to compare clean and poisoned audio data. Figure 10 shows the deviation of corrupted data from clean data norms. This allowed us to understand the effectiveness of the attack.
VII-B1 understanding the similarity of different styles .
Figure 11: Style similarity
Jaccard’s similarity index [94], [95],[96] was used to visualize the similarity between sets of styles ( see Table III) , which we denote by $D_{1},D_{2},D_{3},D_{4},D_{5},$ and $D_{6}$ These similarity ( Figure 11) indices allow us to obtain information about the relationships between different sets of styles, which facilitates an overall understanding of their interconnections.
VII-B2 Speaker verification
VII-B3 Objective :
speaker recognition, on DNN models trained with the TIMIT database poisoned by the DynamicTrigger attack, we apply the speaker verifier(s), ECAPA-TDNN [97](see Figure 12) of HugginFace [^8] HugginFace speaker recognition and Nvidia’s Speakernet [^9] NVIDIA Speech and Translation AI [98] (see Figure 13) to detect irregularities in the final audio data (poisoned) obtained by the DNN model(s) trained on this data in comparison with the clean data.
Figure 12: Speaker verification ECAPA-TDNN.
Figure 13: Speaker verification NeMo Nvidia.
VIII conclusion.
@@ VII-B3 Objective :
This paper presents an efficient dynamic backdoor attack approach for speech recognition by injecting triggers into the auditory speech spectrum. The experimental results illustrate that DynamicTrigger can achieve a sufficiently high ASR while remaining sufficiently stealthy. In addition, DynamicTrigger proves effective in resisting standard defense methods (but is also capable of extending to backdoor attacks via stylistic transformations). Existing backdoor attacks mostly concentrate on classification tasks, while triggers designed for voice command recognition also demonstrate efficacy in other such tasks, including speaker recognition. This article aims to draw the attention of financial services to the adoption of such authentication mechanisms and to encourage them to implement, in addition to these authentication techniques, other means of ensuring that their voice recognition authentication systems are preserved and protected against backdoor attacks.
@@ Acknowledgments.
Ethical statements: Safety and Social Impact.
Backdoor attacks are a serious danger to security; trust, and privacy, thus it’s critical to think through the moral and legal ramifications before developing and implementing backdoor detection methods. Subsequent studies ought to proactively participate in conversations about data security, privacy preservation, and the possible unforeseen effects of implementing such systems.